OpenStudy (aaronandyson):

A car made a run of 390 km in x hours,If the speed had been 4km/hr more , it would have taken 2 hours less for the journey. Find x. @welshfella @sepeario

OpenStudy (aaronandyson):

@Sepeario

OpenStudy (anonymous):

\[x - 2 = \frac{ 390 }{ 4 }\] \[x = 2 + \frac{ 390 }{ 4 } = 99.5\]

OpenStudy (anonymous):

It is a hell of a long trip :D

OpenStudy (aaronandyson):

Whta?

OpenStudy (aaronandyson):

The ans is 15 But how?

OpenStudy (hockeychick23):

ok so you do: (390/x)+4 because the car made the run of 390 km in x hours and you want to know what would happen if you increased the speed 4 km. do you get the first part?

OpenStudy (aaronandyson):

no

OpenStudy (welshfella):

velocity v = distance / time = 390/x v increases by 4 and x is reduced by 2, therefore v + 4 = 390 / (x -2) v = [390 / (x -2] - 4 eliminating v 390/x = [390/x-2) - 4]

OpenStudy (aaronandyson):

Can't I 390/x - 390/x+4 = 2 Do this?

OpenStudy (welshfella):

sorry i got that wrong

OpenStudy (aaronandyson):

390/x - 390/x+2 = 4

OpenStudy (welshfella):

no that not correct

OpenStudy (welshfella):

its